OpenStudy (anonymous):

Lim x approaches 0 5x^2-x/x

OpenStudy (jtvatsim):

The question is to find \[\lim_{x \rightarrow 0} \frac{5x^2 - x}{x}\] right?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

Alright, what have you tried so far? Is there a particular place that's difficult?

OpenStudy (anonymous):

I've tried substitution.

OpenStudy (jtvatsim):

Alright, what did you substitute?

OpenStudy (anonymous):

\[\frac{ 5(3)^{3}-135 }{(3)-3 }\]

OpenStudy (anonymous):

I'm sorry the lim of x going to 3

OpenStudy (anonymous):

not 0

OpenStudy (jtvatsim):

Ah! Big difference... that helps! :)

OpenStudy (jtvatsim):

So, we really are finding \[\lim_{x \rightarrow 3} \frac{5x^2 - x}{x}\]

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, so when you substitute, what did you put for x? was it the equation you typed a bit earlier?

OpenStudy (anonymous):

I substituted x for 3

OpenStudy (jtvatsim):

Good! And what did you get?

OpenStudy (anonymous):

go I got 0/0

OpenStudy (jtvatsim):

How did you get that? The top part of the fraction is 5x^2 - x right?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, so if you just calculate the top part of the fraction what number do you get? Substituting x for 3 in 5x^2 - x

OpenStudy (anonymous):

it's wrong. the substitution for the top is 5x^2-135

OpenStudy (anonymous):

so 5(3)^3-135 =0 and for the bottom would be 3-3 =0

OpenStudy (jtvatsim):

Where did the ^3 cubed come from?

OpenStudy (jtvatsim):

Are you saying that the question was typed in wrong? limit as x approaches 3 of 5x^3 - 135 / x - 3 ?

OpenStudy (jtvatsim):

is that what it should have been? :)

OpenStudy (anonymous):

original problem I posted earlier, yes

OpenStudy (jtvatsim):

OK, OK, that changes things. Then, yes, you are right, we do have a problem with 0/0. So far you have done all the math correctly.

OpenStudy (anonymous):

so after this what would I do? would my answer be (limit does not exist)?

OpenStudy (jtvatsim):

Let me think for just a moment now that I know the real question, and I'll let you know... your suggestion is possible... hang on :)

OpenStudy (jtvatsim):

Aha! Alright, so we need to be a little tricky. Ready for it?

OpenStudy (anonymous):

yes! (:

OpenStudy (jtvatsim):

OK, so, the standard trick when given a limit with a fraction is to try and factor the top part of the fraction so that the bottom part cancels out.

OpenStudy (jtvatsim):

In our case, we want to TRY and factor 5x^3 - 135 so that an (x-3) pops out in the top, which would allow us to cancel the (x-3) in the bottom. This gets rid of the 0/0 problem.

OpenStudy (jtvatsim):

Does that make sense? Anything I can clarify before we actual do it? :)

OpenStudy (anonymous):

no, so far good.

OpenStudy (jtvatsim):

Alright, now do you see anything easy that we can factor out of 5x^3 - 135 ?

OpenStudy (anonymous):

yes.

OpenStudy (jtvatsim):

OK, what would you suggest?

OpenStudy (anonymous):

27?

OpenStudy (jtvatsim):

Well, it might be a bit hard to factor a 27 out of 5x^3 right? Try to ignore that x will be 3 later.

OpenStudy (jtvatsim):

Any other options? Something in common to both 5x^3 and 135 ?

OpenStudy (jtvatsim):

How about 5? Does that seem reasonable?

OpenStudy (anonymous):

Okay, so (x^3-27)

OpenStudy (jtvatsim):

Good, we have 5(x^3 - 27) so far... a little better, but now we have to see if there is a way to factor just x^3 - 27

OpenStudy (jtvatsim):

This is the trick of the problem, so if you'd like I can give you the Algebra to factor it. :)

OpenStudy (anonymous):

please.

OpenStudy (jtvatsim):

lol, OK!

OpenStudy (jtvatsim):

We notice that x^3 - 27 is really just x^3 - 3^3 it is the DIFFERENCE OF PERFECT CUBES

OpenStudy (jtvatsim):

Then, by a trick of Algebra it turns out that x^3 - 3^3 = (x - 3)(x^2 + 3x + 3^2)

OpenStudy (jtvatsim):

That is, x^3 - 27 is the same as (x - 3)(x^2 + 3x + 9)

OpenStudy (jtvatsim):

In fact, this works for all perfect cubes x^3 - n^3 = (x - n)(x^2 + n*x + n^2)

OpenStudy (jtvatsim):

But anyways, we can rewrite our problem as limit as x goes to 3 of 5(x-3)(x^2 + 3x + 9) / (x - 3) now can you solve it?

OpenStudy (jtvatsim):

Of course, if you have any questions, feel free to ask. :)

OpenStudy (anonymous):

thank you, much better.

OpenStudy (jtvatsim):

your welcome! It's a nasty little trick, but it is possible to understand. I just didn't want to bore you with too much Algebra. ;)

OpenStudy (jtvatsim):

Once you cancel out the terms, let me know what you get.